OpenStudy (anonymous):

Solve the following inequality: \[3x^2(x-3)(x+3) ge 0\]

OpenStudy (anonymous):

\[3x^2(x-3)(x+3) \ge 0\]

OpenStudy (lgbasallote):

dont you just equate each factor to 0? \[3x^2 \ge 0\] \[x-3 \ge 0\] \[x+3 \ge 0\] solve for x in each one

OpenStudy (anonymous):

I thought your not allow to separate it?

OpenStudy (anonymous):

Oh wait, never mind, that's set it equal to zero!... okay, thanks @lgbasallote

OpenStudy (lgbasallote):

haha yes :)

OpenStudy (anonymous):

answer is \[\left| x \right| \ge 3\]

OpenStudy (anonymous):

u were right wer not allow to separate it